# Boosting MLLM Spatial Reasoning with Geometrically Referenced 3D Scene Representations

### 09 Mar 2026 v1

@@ Abstract
While Multimodal Large Language Models (MLLMs) have achieved remarkable success in 2D visual understanding, their ability to reason about 3D space remains limited. To address this gap, we introduce geometrically referenced 3D scene representations (GR3D). Given a set of input images, GR3D annotates objects in the images with unique IDs and encodes their 3D geometric attributes as textual references indexed by these IDs. This representation enables MLLMs to interpret 3D cues using their advanced language-based skills in mathematical reasoning, while concurrently analyzing 2D visual features in a tightly coupled way. We present a simple yet effective approach based on GR3D, which requires no additional training and is readily applicable to different MLLMs. Implemented in a zero-shot setting, our approach boosts GPT-5’s performance on VSI-Bench by 8% overall and more than 11% on tasks that rely heavily on spatial layout understanding. Qualitative studies further demonstrate that GR3D empowers MLLMs to perform complex spatial reasoning with highly sparse input views.
@@ 1 Introduction
Figure 1: An overview of GR3D framework. Given a collection of images, our method reconstructs 3D scenes, extracts object-level geometric attributes, and transform them into a GR3D representation, which consists of annotated images and textual references of geometric attributes. Such paired text and images are provided to a MLLM to perform spatial reasoning tasks.
An overall framework of leveraging GR3D for spatial reasoning is illustrated in Fig. 1. GR3D generation builds on recent advances in neural 3D reconstruction. Methods such as DUSt3R [33] and VGGT [31] can reconstruct 3D scenes from uncalibrated, unposed images with high accuracy and consistency. From reconstructed scenes, we derive object geometric attributes (e.g., bounding boxes, primitive types, and shape parameters) and convert them into structured textual references indexed by object IDs. Using estimated camera parameters, we project object centers into images and annotate projected locations with corresponding objects IDs. The annotated images replace the original images as visual input for MLLMs, while textual references are combined with task-specific instructions to form text prompts. Our method does not require training and can work along with various MLLMs.
We evaluate our approach using the Visual-Spatial Intelligence Benchmark (VSI-Bench) [39], which contains over 5,000 question-answer pairs derived from 288 indoor videos. The questions cover different types of spatial understanding tasks. We implement our method in a zero-shot manner. We neither fine-tune on any data that overlap with this benchmark dataset nor construct similar question-answer pairs for training. Nevertheless, our method achieves state-of-the-art benchmarking performance, demonstrating strong generalization and effectiveness. Ablation studies further confirm the critical role of GR3D in achieving this performance.
@@ 3 Method
Given a set of multi-view images, we employ neural 3D reconstruction methods such as DUSt3R [33] to recover 3D scenes and camera poses. Unlike traditional Structure-from-Motion (SfM) pipelines that rely on handcrafted keypoint matching, such models use dense feature correlation and neural optimization, yielding reconstructions that are highly robust under textureless regions, occlusions, and wide baseline conditions. The output consists of a dense depth map $D$ for each input image along with the corresponding Camera intrinsics $K$ and extrinsics $(R,t)$. A 3D point cloud in a global coordinate system can be derived by unprojecting pixels with their depth values. Since reconstructions are defined only up to an unknown scale factor, the metric scale can be recovered either from auxiliary depth sensor measurements, when available, or from reference objects of known dimensions.
@@ 3.2 Language-Based Geometric Reasoning
Another advantage of using languages for geometric reasoning is the enhanced explainability, which is mostly absent in smaller fine-tuned models. Modern MLLMs naturally generate chain-of-thought reasoning, in which intermediate steps are explicitly articulated. The step-by-step calculation and reasoning not only reveal how a final answer is derived but also facilitate error analysis and model refinement. By examining intermediate steps, it becomes possible to identify systematic mistakes, assess model reliability, and design targeted strategies [45, 36] to reduce errors and enhance performance.
@@ 3.3 Text-Image Association
When provided with both images and geometric descriptions, MLLMs often struggle to associate the two modalities, particularly in complex scenes containing many objects. This challenge is evident in our experiments, as discussed in Sec. 4.1.4. To address this issue, we annotate objects on images with their IDs, which establishes a clear correspondence between visual and geometric information. Given an object with geometric attributes, we project its object center $C$ to the image by first transforming it into camera canonical coordinates:
@@
During object center projection, no occlusion check is performed. As a result, occluded objects are also marked, leading to incorrect association. This issue is illustrated in the left image in Fig. 2. While one solution is to render full geometric shapes for an occlusion check, it significantly increases computational cost. Instead, we exploit the depth maps $D$ that are output from 3D reconstruction models. Specifically, we do not annotate an object if $z$ in Eq. 1 is smaller than $D_{(i,j)}$, i.e., the object center is farther than the depth at the same location. As can be seen in Fig. 2, most annotations corresponding to occluded objects are removed.
@@ 4 Experiments
In this section, we quantitively evaluate our proposed approach and compare its performance with state-of-the-art models. We also demonstrate the effectiveness of our approach in scenarios with highly sparse input views.
4.1 Quantitative Evaluation
4.1.1 Datasets
We evaluate our approach on the Visual-Spatial Intelligence Benchmark (VSI-Bench) [39], a large-scale benchmark designed to measure visual spatial intelligence. The dataset covers a broad spectrum of spatial understanding tasks. These include configurational tasks including object counting, relative distance, relative direction, and route planning, measurement estimation tasks including object size, room size, and absolute distance, and a spatiotemporal task of determining the order of appearance. We select VSI-Bench for evaluation because it is, to our knowledge, the largest and most comprehensive dataset for evaluating spatial reasoning on complex scenes. Other existing benchmarks are less extensive [34], combine single-view and multi-view settings [16, 35], or focus on object-centric reasoning [40].
4.1.2 Implementation Details
VSI-Bench provides videos of indoor environments. For reconstruction, we apply VGGT [31] to subsampled frames. We also run Mask2Former [8] trained on ADE20K to obtain semantic segmentation for each frame. The reconstructed point clouds are rotated to be axis-aligned using the up direction estimated from floor points and the dominant direction of horizontal room boundaries. We recover a global scaling factor from the ratio between the typical real-world heights and reconstructed heights of selected objects (e.g., ceilings, countertops, and desks). Although approximate, this estimate enables measurement-related tasks.
@@ 4.1.2 Implementation Details
The GR3D input consists of annotated images and textual references inserted into a prompt template (see Fig. 3). The prompt explains the structure of the representation. We observe that it is important to specify the handedness and the upright orientation of the coordinate system, as this allows MLLMs to correctly apply orientation tests to solve direction-related questions. To reduce token usage, we provide the pre-computed room size derived from room boundary polygons, while the polygons can be included for tasks that require reasoning over room geometry.
Figure 3: Prompt template for VSI-Bench evaluation.
4.1.3 Evaluation Results
Evaluation results on VSI-Bench are summarized in Tab. 1. For comparison, we include the benchmark scores of GPT-4o and Gemini-1.5 Pro, reported in [39], as well as the scores of InternVL2 [6], a leading open-source mid-size model on this benchmark. VG LLM [43] is among the few models fine-tuned with 3D scene data and evaluated on VSI-Bench. In addition, we evaluate the recently released GPT-5111We set reasoning effort to low, which provides a favorable balance between accuracy and efficiency. on this benchmark. To assess the effectiveness of our method, we measure the performance of incorporating GR3D with InternVL2, GPT-4o and GPT-5222We follow the setting in [39], using 32 frames for InternVL2 and 16 frames for the GPT models..
@@ 4.1.3 Evaluation Results
Among all MLLMs, GPT-5 achieves the highest overall score. While the fine-tuned VG LLM attains a comparable performance, it should be noted that its training data include videos overlapping with VSI-Bench and questions similar to those in the benchmark. When integrated with GR3D, all three models exhibit performance gains, where models with stronger reasoning capabilities show greater improvements. In particular, GPT-5 achieves a 8% increase, setting a new state of the art. Notably, for configurational tasks including object count, relative distance, relative direction, and route planning, GR3D yields substantial improvements of 14%, 15%, 20%, and 11%, respectively. This validates that our approach represents 3D spatial information in a form that can be effectively utilized by MLLMs. The experimental results demonstrate that our approach can function as a general plug-in mechanism for enhancing spatial reasoning in MLLMs. As MLLMs continue to advance rapidly, our approach can immediately leverage these advancements without fine-tuning or retraining.
While our approach achieves strong overall performance, further analysis reveals that a large portion of the errors stem from inaccurate geometric attributes. In particular, our current implementation mainly relies on voxel connectivity to separate individual object instances, which struggles in cluttered scenes and thus reduces the accuracy of object counting and size estimation. More advanced 3D instance segmentation models could lead to improvements. In addition, when point clouds are highly noisy, the current object detection method often overestimates bounding box dimensions. This issue can be mitigated through more robust 3D semantic segmentation or denoising techniques.
4.1.4 Ablation Study
In this section, we conduct ablation studies. For a fair comparison, all experiments are performed using GPT-5. Results are summarized in Tab. 2. For brevity, we present scores of three configurational tasks and the average score of all tasks.
@@ 4.1.4 Ablation Study
Method Rel.Dist. Rel.Dir. RoutePlan Avg.
GR3D 57.6 62.2 61.2 58.7
w/o annotation 52.8 51.3 52.0 50.3
Camera param. 40.3 42.6 45.4 48.8
Scene desc. 48.6 51.9 51.0 53.4
Table 2: Ablation study. We evaluate 1) removing object annotations in images, 2) providing camera parameters to implicitly associate images with object geometric information, and 3) converting images into scene descriptions with object ID citations. For brevity, we show the scores of three configurational tasks and the overall average.
4.2 Sparse View Spatial Reasoning
While spatial reasoning benchmarks often assume access to densely sampled video frames, in many real-world scenarios, only a limited number of input views are available. Examples include property listing photos on real estate, hotel booking, and short-term rental platforms, which usually capture only selected angles of rooms. Similarly, surveillance systems often rely on fixed-angle cameras that offer sparse coverage of a scene. In such cases, reasoning about spatial layout and object relationships is highly valuable yet inherently challenging.
@@ 4.2 Sparse View Spatial Reasoning
In Fig. 4, we present example results from two scenes in VSI-Bench, each of which is captured by only four images subsampled from input videos. Four questions are posed, which require complex multi-step reasoning and an understanding of spatial context from an egocentric perspective. The questions are designed such that human can infer the answers from the images. We apply the same approach described in Sec. 4.1.2. For comparison, we also show the responses from GPT-5 alone, which receives the questions and unannotated images as input. Based on the results, we make the following observations:
@@
  • GR3D allows GPT-5 to combine 3D geometric reasoning with its strong 2D visual understanding, successfully identifying uncommon objects such as the foam roller.
These results show that GR3D enables GPT-5 to accurately infer object arrangements and answer spatial queries even with very sparse input views. This suggests strong potential for deploying our approach where dense visual data collection is infeasible.

### 26 Apr 2026 v2

@@ Abstract
While Multimodal Large Language Models (MLLMs) have achieved remarkable success in 2D visual understanding, their ability to reason about 3D space remains limited. To address this gap, we introduce geometrically referenced 3D scene representations (GR3D). Given a set of input images, GR3D annotates objects in the images with unique IDs and encodes their 3D geometric attributes as textual references indexed by these IDs. This representation enables MLLMs to interpret 3D cues using their advanced language-based skills in mathematical reasoning, while concurrently analyzing 2D visual features in a tightly coupled way. We present a simple yet effective approach based on GR3D, which requires no additional training and is readily applicable to different MLLMs. Implemented in a zero-shot setting, our approach yields substantial improvements on challenging spatial reasoning benchmarks, boosting GPT-5 performance by 9% on VSI-Bench and 12% on MindCube. Qualitative studies further demonstrate that GR3D empowers MLLMs to perform complex spatial reasoning with highly sparse input views.
@@ 1 Introduction
Figure 1: An overview of GR3D framework. Given a collection of images, our method reconstructs 3D scenes, extracts object-level geometric attributes, and transforms them into a GR3D representation, which consists of annotated images and textual references of geometric attributes. Such paired text and images are provided to a MLLM to perform spatial reasoning tasks.
An overall framework of leveraging GR3D for spatial reasoning is illustrated in Fig. 1. GR3D generation builds on recent advances in neural 3D reconstruction. Methods such as VGGT [33] and $\pi^{3}$ [36] can reconstruct 3D scenes from uncalibrated, unposed images with high accuracy and consistency. From reconstructed scenes, we derive object geometric attributes (e.g., bounding boxes, primitive types, and shape parameters) and convert them into structured textual references indexed by object IDs. Using estimated camera parameters, we project object centers into images and annotate projected locations with corresponding object IDs. The annotated images replace the original images as visual input for MLLMs, while textual references are combined with task-specific instructions to form text prompts. Our method does not require training and can work along with various MLLMs.
We evaluate our approach on two challenging spatial reasoning benchmarks, the Visual-Spatial Intelligence Benchmark (VSI-Bench) [39] and MindCube [40]. We implement our method in a zero-shot manner. We neither fine-tune on any data that overlap with this benchmark dataset nor construct similar question-answer pairs for training. Nevertheless, our method achieves state-of-the-art benchmarking performance, demonstrating strong generalization and effectiveness. Ablation studies further confirm the critical role of GR3D in achieving this performance.
@@ 3 Method
Given a set of multi-view images, we employ neural 3D reconstruction models to recover 3D scenes and camera poses. Unlike traditional Structure-from-Motion (SfM) pipelines that rely on handcrafted keypoint matching, such models use dense feature correlation and neural optimization, yielding reconstructions that are highly robust under textureless regions, occlusions, and wide baseline conditions. The output consists of a dense depth map $D$ for each input image along with the corresponding Camera intrinsics $K$ and extrinsics $(R,t)$. A 3D point cloud in a global coordinate system can be derived by unprojecting pixels with their depth values.
@@ 3.2 Language-Based Geometric Reasoning
Another advantage of using languages for geometric reasoning is the enhanced explainability, which is mostly absent in smaller fine-tuned models. Modern MLLMs can generate chain-of-thought reasoning, in which intermediate steps are explicitly articulated. The step-by-step calculation and reasoning not only reveal how a final answer is derived but also facilitate error analysis and model refinement. By examining intermediate steps, it becomes possible to identify systematic mistakes, assess model reliability, and design targeted strategies [45, 35] to reduce errors and enhance performance.
@@ 3.3 Text-Image Association
When provided with both images and geometric descriptions, MLLMs often struggle to associate the two modalities, particularly in complex scenes containing many objects. This challenge is evident in our experiments, as discussed in Sec. 4.4. To address this issue, we annotate objects on images with their IDs, which establishes a clear correspondence between visual and geometric information. Given an object with geometric attributes, we project its object center $C$ to the image by first transforming it into camera canonical coordinates:
@@
During object center projection, no occlusion check is performed. As a result, occluded objects are also marked, leading to incorrect association. This issue is illustrated in the left image in Fig. 2. While one solution is to render full geometric shapes for an occlusion check, it significantly increases computational cost. Instead, we exploit the depth maps $D$ that are output from 3D reconstruction models. Specifically, we do not annotate an object if $z$ in Eq. 1 is larger than $D_{(i,j)}$, i.e., the object center is farther than the depth at the same location. As can be seen in Fig. 2, most annotations corresponding to occluded objects are removed.
@@ 4 Experiments
4.1 Benchmarks
We evaluate our approach on VSI-Bench [39], and MindCube [40]. VSI-Bench contains over 5,000 questions derived from 288 in-door videos. This benchmark assesses a broad spectrum of spatial understanding capabilities, including configurational tasks (object counting, relative distance, relative direction, and route planning), measurement estimation tasks (object size, room size, and absolute distance), and a spatiotemporal task of determining the order of appearance. MindCube focuses on spatial reasoning from limited input views. It contains questions on object spatial relationships, where each question is based on a scene captured by 2 to 4 views, with various camera movement types.
4.2 Implementation Details
We adopt the $\pi^{3}$ model [36] for reconstruction, as it achieves high accuracy and can predict metric scale. For video input, It is applied to subsampled frames. We also run Mask2Former [8] trained on ADE20K to obtain semantic segmentation for each input image. The reconstructed point clouds are rotated to be axis-aligned using the up direction estimated from floor points and the dominant direction of horizontal room boundaries.
@@ 4.2 Implementation Details
The GR3D input consists of annotated images and textual references inserted into a prompt template. Fig. 3 shows the template used in the experiments, which explains the structure of the representation. We observe that it is important to specify the handedness and the upright orientation of the coordinate system, as this allows MLLMs to correctly apply orientation tests to solve direction-related questions. To reduce token usage, we provide the pre-computed area of room boundary polygons for room size questions.
Figure 3: Prompt template used in evaluations.
4.3 Evaluation Results
Evaluation results on VSI-Bench are summarized in Tab. 1. For comparison, we include the benchmark scores of GPT-4o and Gemini-1.5 Pro, reported in [39], as well as the scores of InternVL2 [6], a leading open-source mid-size model on this benchmark. VG LLM [43] is among the few models fine-tuned with 3D scene data and evaluated on VSI-Bench. In addition, we evaluate GPT-5 on this benchmark  [^1] We set reasoning effort to low, which provides a favorable balance between accuracy and efficiency.. To assess the effectiveness of our method, we measure the performance of incorporating GR3D with InternVL2, GPT-4o and GPT-5  [^2] We follow the setting in [39], using 32 frames for InternVL2 and 16 frames for the GPT models..
@@ 4.3 Evaluation Results
Among all MLLMs, GPT-5 achieves the highest overall score. While the fine-tuned VG LLM attains a comparable performance, it should be noted that its training data include videos overlapping with VSI-Bench and questions similar to those in the benchmark. When integrated with GR3D, all three models exhibit performance gains, where models with stronger reasoning capabilities show greater improvements. In particular, GPT-5 achieves a 9% increase, setting a new state of the art. This validates that our approach represents 3D spatial information in a form that can be effectively utilized by MLLMs. The improvements across all three models demonstrate that our approach can function as a general plug-in mechanism for enhancing spatial reasoning in MLLMs. As MLLMs continue to advance rapidly, our approach can immediately leverage these advancements without fine-tuning or retraining.
Tab. 2 presents the MindCube results. We include models with leading performance reported in  [40]. Consistent with the trends observed on VSI-Bench, GR3D yields substantial performance improvements, boosting the accuracy of GPT-5 by 12%. To our knowledge, this result represents the highest accuracy achieved by zero-shot methods on MindCube, highlighting the effectiveness of our representation in scenarios with sparse input views. Overall, the consistent gains across both VSI-Bench and MindCube demonstrate the generality of our approach and its ability to support diverse forms of 3D spatial reasoning without task-specific adaptation.
While our approach achieves strong overall performance, further analysis reveals that a large portion of the errors stem from inaccurate geometric attributes. In particular, our current implementation mainly relies on voxel connectivity to separate individual object instances, which struggles in cluttered scenes and thus reduces the accuracy of object counting and size estimation. More advanced 3D instance segmentation models could lead to improvements. In addition, when point clouds are highly noisy, the current object detection method tends to either miss objects or overestimate their dimensions. This issue can be mitigated through more robust 3D semantic segmentation or denoising techniques.
Method Name Rotation Among Around Overall
LLaVA-Onevision-7B [18] 36.5 48.4 44.1 47.4
DeepSeek-VL2-Small [22] 37.0 50.4 26.9 47.6
GPT-5-2025-08-07 [14] 87.5 39.3 66.4 55.0
GR3D + GPT-5-2025-08-07 88.5 55.7 76.0 66.8
Table 2: MindCube evaluation results. Rotation, among, and around represent subsets with different camera placements. GR3D combined with GPT-5 achieves the highest accuracy. The best result for each subset across all models is indicated in bold.
4.4 Ablation Study
In this section, we conduct ablation studies. We use VSI-Bench and GPT-5 for all experiments. Results are summarized in Tab. 3. For brevity, we present scores of two configurational tasks and the average score of all tasks.
@@ 4.4 Ablation Study
Impact of reconstruction models. To assess the impact of the reconstruction component, we replace $\pi^{3}$ with VGGT. As VGGT does not predict metric scales, we estimate a scaling factor from the ratio between the typical real-world heights and reconstructed heights of selected objects (e.g., ceilings, countertops, and desks). As can be seen, $\pi^{3}$ performs slightly better than VGGT, due to its improved reconstruction accuracy. We expect that continued advances in reconstruction models will further improve overall performance.
@@
Method Rel. Dir. Route Plan Avg.
GR3D 69.5 58.7 59.8
No object annotation 51.3 52.0 50.3
VGGT reconstruction 62.2 56.7 58.3
Camera parameters 42.6 45.4 48.8
Scene descriptions 51.9 51.0 53.4
Table 3: Ablation study. We evaluate 1) removing object annotations in images, 2) replacing $\pi^{3}$ with VGGT, 3) providing camera parameters to implicitly associate images with object geometric information, and 4) converting images into scene descriptions with object ID citations. For brevity, we show the scores of two configurational tasks and the overall average.
4.5 Qualitative Examples
@@
To provide further insight on how GR3D improves spatial reasoning, we present qualitative examples that highlight its impact on model behavior in challenging cases. In Fig. 4, we present example results from two scenes in VSI-Bench, each of which is captured by only four images subsampled from input videos. Four questions are posed, which require complex multi-step reasoning and an understanding of spatial context from an egocentric perspective. The questions are designed such that human can infer the answers from the images. We apply the same approach described in Sec. 4.2. For comparison, we also show the responses from GPT-5 alone, which receives the questions and unannotated images as input. Based on the results, we make the following observations:
@@
  • GR3D allows GPT-5 to perform 3D geometric reasoning while simultaneously leveraging its strong 2D visual understanding. For example, it can identify uncommon objects such as the foam roller.